# Predicting Multi-Agent Specialization via Task Parallelizability

### v1

@@ A.3 Training Procedures
For each trial in our MARL experiments, we trained two independent agents to play Overcooked collaboratively through self-play. Our implementation builds upon the PureJaxRL framework for Proximal Policy Optimization (Schulman et al., 2017; Lu et al., 2022). Each agent was implemented using an soft Actor-Critic network (Haarnoja et al., 2018), in which both the actor and critic networks were composed of two fully-connected layers with 64 hidden units each. We trained each agent using Proximal Policy Optimization with an Adam optimizer and a learning rate of $2.5\times 10^{-3}$ for $1\times 10^{7}$ timesteps (Zhang, 2018). At each timestep, the agents observed the state of the current environment as a $\text{width}\times\text{height}\times 27$-dimensional grid. The first two dimensions corresponded to the spatial coordinates of the grid, while the 27-dimensional vector encoded the state of each grid cell, indicating whether it contained an agent, onion, tomato, bowl, or workstation. The policy produced a categorical distribution over discrete actions.
@@
To accelerate learning in the sparse-reward environments, we applied reward shaping to augment raw rewards with an auxiliary-shaped signal. During reward shaping, each agent received rewards for performing specific actions aligned with the given recipe’s requirements: adding onions or tomatoes to the pot when an additional ingredient is needed, cooking the pot once it contains the correct quantity of ingredients, retrieving the soup, and finally serving it. The shaping reward $r_{s}$ was annealed over a predefined horizon ($2.5\times 10^{6}$) using a linear schedule.
A.4 Experiment 1: Significance Tests
@@
Number of Pots and Workstations. First, we found a significant effect of the number of pots, but not the total number of workstations, on specialization. Agents tended to specialize more with one pot ($M=0.77,SD=0.30$) than two ($M=0.57,SD=0.41$) but not when presented with more of other workstations (e.g., bowls, serving stations). A Welch’s t-test between layouts with one pot versus two revealed a significant difference in JSD, $t(2842.72)=14.98,p=7.23\times 10^{-49}$. Conversely, a one-way ANOVA showed a non-significant difference in JSD across the full range of possible workstation combinations ($F(5,3094)=0.82,p=0.54$).
@@
Spatial Bottlenecks. Next, we found a significant effect of spatial bottlenecks on specialization. Agents specialized less in layouts with fewer bottlenecks ($M=0.61,SD=0.38$) than in those with medium ($M=0.69,SD=0.37$) or high bottlenecks ($M=0.86,SD=0.27$). Using edge betweenness centrality, we quantified the average extent of spatial bottlenecks across the entire layout. Betweenness values were divided into three equal-width bins: Low $[2.110,3.194)$, Medium $(3.195,4.277]$, and High $(4.277,5.360]$. Smaller centrality values represented a fewer number of nodes that were critical to many paths in the layout graph, corresponding to a smaller number of spatial bottlenecks. Larger centrality values represented more critical nodes, corresponding to greater spatial bottlenecks. A one-way ANOVA showed a significant difference in JSD across bins ($F(2,2097)=34.10,p=2.26\times 10^{-15}$). Pairwise comparisons using Tukey’s HSD showed that JSD was significantly higher with medium centrality than low ($p=3.18\times 10^{-7}$), high centrality than medium ($p=1.07\times 10^{-12}$), and high centrality than low ($p=3.82\times 10^{-20}$).
@@
Recipe Complexity. Finally, we found a significant effect of recipe complexity (the number of ingredients required to make a soup) on JSD. Agents tended to specialize less when recipes had one ingredient ($M=0.54,SD=0.47$) than with two ($M=0.73,SD=0.35$) or three ($M=0.67,SD=0.32$). A one-way ANOVA showed a significant difference in JSD across recipes ($F(2,3097)=57.22,p=3.95\times 10^{-25}$). Pairwise comparisons using Tukey’s HSD revealed that JSD was significantly higher for recipes with two ingredients compared to one ($p=4.46\times 10^{-19}$) and recipes with three ingredients compared to one ($p=4.37\times 10^{-10}$), but significantly lower for recipes with three ingredients compared to two ($p=7.77\times 10^{-6}$).
A.5 Experiment 1: Evaluation
@@
We also fit a logistic regression model where the binary outcome variable was low versus high specialization (JSD $\geq 0.5$) and the predictor variable was task parallelizability. To address the class imbalance in the JSD distribution, we used class-weight balancing to automatically adjust class weights inversely proportional to their frequencies in the training data.
Figure 6: Overcooked Spatial Layouts in Experiment 1.
Figure 7: Effect of grid size on observed specialization (JSD) in Experiment 1. As the size of the grid increased, agends tended to specialize more (even when presented with the same number of workstations and bottlenecks). W
Spatial Layout Mean JSD Mean Reward
@@
Table 1: Average specialization (JSD) observed in different spatial layouts in Experiment 1.
Workstations Mean JSD Std JSD
@@
Table 2: Mean and standard deviation of observed specialization (JSD) for varying numbers of workstations in Experiment 1.
Recipes Mean JSD Std JSD
@@
Table 3: Mean and standard deviation observed specialization (JSD) for different recipes in Experiment 1.
@@
Table 5: Correlations between predicted task parallelizability ($S$) and observed specialization (JSD) across varying spatial layouts in Experiment 1. The significance of this relationship varied across layouts, especially those with different grid sizes, prompting the controlled design of Experiments 2 and 3.
@@
Figure 8: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific spatial layout.
@@
Figure 9: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution for a specific number of workstations.
@@
Figure 10: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific recipe.
@@
Figure 11: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific layout. The distributions of strategies appear mostly consistent across different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
@@
Figure 12: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by the number of workstations. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
@@
Figure 13: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific recipe. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
A.6 Experiment 2: Results
Number of Pots Mean JSD Mean Reward
1 0.676053 14.861111
2 0.280481 18.867647
Table 6: Average specialization (JSD) with varying numbers of pots in Experiment 2.
Spatial Layout Mean JSD Mean Reward
Open 0.3320525 18.7083330
Horizontal Divider 0.332101 18.0416668
Vertical Divider 0.318273 18.0000000
Vertical Space 0.404964 18.0833335
Long Divider 0.752011 12.2083334
Two Dividers 0.2988275 17.6666666
Table 7: Average specialization (JSD) in various spatial layouts in Experiment 2.
A.7 Experiment 3: Design
Figure 14: Three layouts from Experiment 3 are shown: (a) the smallest layout, (b) the largest layout with increased exploration, and (c) the largest layout with consistent exploration. Across these configurations, the optimal policy remains constant. The plots depict how the JSD and reward vary with increasing layout size under two different conditions. In the case of the three-onion soup, the reward decreases as the layout size grows. The smallest layout (layout 1) corresponds to the minimum JSD and the maximum reward.
In Experiment 3, we examined the influence of various divider configurations on agents specialization. Four skeleton layouts of identical size were tested: (1) a single vertical divider, (2) a double vertical divider, (3) a horizontal divider, and (4) no divider. The focal point was to investigate how these structural variations affect the convergence of agent strategies. To evaluate the impact of state space size on optimal policy convergence and in cases of suboptimal convergence, whether agents developed more specialist or generalist strategies - we conducted a controlled experiment comprised of 280 runs. The experiment systematically varied seven layout sizes, utilizing 10 random seeds per layout, two recipe configurations (1-onion soup and 3-onion soup), and two experimental conditions (constant vs. non-constant exploration). Critically, the optimal policy remained consistent across all runs; only empty space that the agents could not access was added to increase the size of the state space. Figure 14 illustrates three of these layouts, demonstrating how variations in layout size were implemented, while keeping the optimal policy consistent
A.8 Experiment 3: Results
For the three-onion soup task, the majority of run converged to a JSD value fo approximately 0.6 - achieved in $93.8\%$ of runs with a single divider, $87.5\%$ of runs with a double divider, $87.5\%$ in the no-divider room layout, and $93.8\%$ in the counter circuit layout. Notably, a JSD value of 0 corresponds to the optimal strategy - however, agents rarely reached this level of convergence. For the one-onion soup tasks, full specialization was observed in most runs across the four layouts; $100\%$ with a single divider, $81.3\%$ with a double divider, and $81.3\%$ with no-divider. In stark contrast, only $46.3\%$ of runs in the counter circuit layout converged to a specialist strategy ($JSD>0.5$), while the remaining $53.4\%$ of runs resulted in generalist behavior.
The differences in emergent strategies between layouts is likely due to the distinct effects of the dividers on agent behavior. In layouts with vertical dividers, agents are more inclined to specialize, as the dividers channel them toward the same region where they cannot work in parallel. Conversely, the counter circuit layout encourages generalist strategies by focusing them to either occupy separate regions or follow each other, reducing the likelihood of specialization.
The effect of layouts size varied between the two recipes. For the simpler 1-onion soup recipe, there was no significant relationship between layout size and either reward JSD. Under non-constant exploration, JSD increased significantly with layout size ($r=0.7924,p=0.0336$) indicating greater policy specialisation in large layouts, but reward showed no reliable trend ($r=-0.2512,p=0.5869$). Pairwise T-tests confirmed no significant difference between the smallest and largest environments for JSD ($p=0.755$ or reward $p=0.210$. Under constant exploration, the 1-onion soup was unaffected by layout size, with no substantial correlation for JSD ($r=-0.5850,p=0.1677$) or reward ($r=0.0609,p=0.8969$).
In contrast, the 3-onion soup recipe was sensitive to increasing layout size. Under non-constant exploration, as layout size increased, JSD rose significantly ($r=0.7549,p=0.0498$) while reward decreased ($r=-0.8556,p=0.0140$), indicating that agents struggle to perform well in large environments and develop more specialised policies. Pairwise T-tests further outlined significant differences between the smallest and largest layouts for JSD ($p=0.008$ and reward $p<10^{-4}$. Even under constant exploration, the 3-onion soup recipe showed a decline in reward as layout size increased ($r=-0.4308,p=0.0020$), although JSD was not significantly affected.
These findings demonstrate that increasing both layout size and the number of subtasks, thus expanding the exploration space, drives agents toward suboptimal and overly specialized policies. Even when exploration remains constant, the addition of extraneous state dimensions in larger layouts prevents the RL policy’s ability to converge effectively to the optimal solution. This underscores a key challenge for future work: MARL policies struggle to consistently identify optimal generalist solutions in larger layouts, even when task structure, workstations, and exploration dynamics remain unchanged.
Layout Experiment Recipe Mean JSD Std JSD Count Mean Reward Std Reward Reward Count
@@ A.8 Experiment 3: Results
Table 8: Detailed results from Experiment 3. Each row represents a unique combination of spatial layout, experiment, and recipe, showing the mean, standard deviation (std), and count (number of runs) for both specialization (JSD) and reward metrics. Experiment A corresponds to expanding exploration, while Experiment B denotes constant exploration. The JSD count is less than 10 for certain permutations because the policy failed to converge, resulting in the agents taking no meaningful actions.
@@
Table 9: Correlations from Experiment 3. Correlations between specialization (JSD) and reward for one-onion and three-onion Soup recipes under expanding and constant exploration conditions. The table reports correlations and p-values ($p$) for JSD and reward in relation to layout size.
@@
Table 10: Results from Experiment 3. Pairwise T-tests comparing the smallest (Layout 1) layout with the largest layout (Layout 67).

### v2

@@ Appendix E Training Procedures
For each trial, we trained two independent agents to play SMAC, MPE, or Overcooked collaboratively through self-play. Our implementation builds upon the PureJaxRL and JaxMARL frameworks for Proximal Policy Optimization [42, 43, 41]. Each SMAC and MPE agent was implemented using a recurrent actor-critic network. Observations were embedded using a fully connected ReLU layer, followed by a gated recurrent unit (GRU) to model temporal dependencies. Each Overcooked-AI agent was implemented using an soft actor-critic network [65], in which both the actor and critic networks were composed of two fully-connected layers with 64 hidden units each. We trained each agent using Independent PPO with an Adam optimizer and a learning rate of $2.5\times 10^{-3}$ for $1\times 10^{7}$ timesteps [66].
For SMAC, at each timestep, each agent received an observation vector containing local features about itself and other units within its field of view. This included its own health, position, unit type, and weapon cooldown, as well as relative information about nearby allies and enemies (e.g., health, positions, last movement direction, and attack status). The global state included information about all units on the battlefield and was only used during training. The policy produced a categorical distribution over a discrete set of actions, including movement, attack actions, and wait.
For MPE, at each timestep, each agent received a continuous observation vector that encoded its own velocity and position, the relative positions of other agents, and the relative positions of static landmarks. Agents operated in a continuous 2D space. The policy output a categorical distribution over these actions, with masking used when necessary (e.g., to prevent collisions or enforce environment constraints).
For Overcooked-AI, at each timestep, the agents observed the state of the current environment as a $\text{width}\times\text{height}\times 27$-dimensional grid. The first two dimensions corresponded to the spatial coordinates of the grid, while the 27-dimensional vector encoded the state of each grid cell, indicating whether it contained an agent, onion, tomato, bowl, or workstation. The policy produced a categorical distribution over discrete actions.
@@
To accelerate learning in the sparse-reward environment of Overcooked-AI, we applied reward shaping to augment raw rewards with an auxiliary-shaped signal. During reward shaping, each agent received rewards for performing specific actions aligned with the given recipe’s requirements: adding onions or tomatoes to the pot when an additional ingredient is needed, cooking the pot once it contains the correct quantity of ingredients, retrieving the soup, and finally serving it. The shaping reward $r_{s}$ was annealed over a predefined horizon ($2.5\times 10^{6}$) using a linear schedule.
Compute
All experiments were run on A100 GPUs, totaling 1,569 GPU-hours across 17,813 runs. Experiment 1 used 93 GPU-hours (160 runs on MPE and SMAC). Experiment 2 used 120 GPU-hours (1,440 runs), Experiment 3 used 1,333 GPU-hours (16,000 runs), and Experiment 4 used 23 GPU-hours (280 runs).
Appendix F Experiment 2: Results
Number of Pots Mean SI Mean Reward
1 0.676053 14.861111
2 0.280481 18.867647
Table 1: Average specialization (SI) with varying numbers of pots in Experiment 2.
Spatial Layout Mean SI Mean Reward
Open 0.3320525 18.7083330
Horizontal Divider 0.332101 18.0416668
Vertical Divider 0.318273 18.0000000
Vertical Space 0.404964 18.0833335
Long Divider 0.752011 12.2083334
Two Dividers 0.2988275 17.6666666
Table 2: Average specialization (SI) in various spatial layouts in Experiment 2.
Appendix G Experiment 3: Significance Tests
@@
Number of Pots and Workstations. First, we found a significant effect of the number of pots, but not the total number of workstations, on specialization. Agents tended to specialize more with one pot ($M=0.77,SD=0.30$) than two ($M=0.57,SD=0.41$) but not when presented with more of other workstations (e.g., bowls, serving stations). A Welch’s t-test between layouts with one pot versus two revealed a significant difference in SI, $t(2842.72)=14.98,p=7.23\times 10^{-49}$. Conversely, a one-way ANOVA showed a non-significant difference in SI across the full range of possible workstation combinations ($F(5,3094)=0.82,p=0.54$).
@@
Spatial Bottlenecks. Next, we found a significant effect of spatial bottlenecks on specialization. Agents specialized less in layouts with fewer bottlenecks ($M=0.61,SD=0.38$) than in those with medium ($M=0.69,SD=0.37$) or high bottlenecks ($M=0.86,SD=0.27$). Using edge betweenness centrality, we quantified the average extent of spatial bottlenecks across the entire layout. Betweenness values were divided into three equal-width bins: Low $[2.110,3.194)$, Medium $(3.195,4.277]$, and High $(4.277,5.360]$. Smaller centrality values represented a fewer number of nodes that were critical to many paths in the layout graph, corresponding to a smaller number of spatial bottlenecks. Larger centrality values represented more critical nodes, corresponding to greater spatial bottlenecks. A one-way ANOVA showed a significant difference in SI across bins ($F(2,2097)=34.10,p=2.26\times 10^{-15}$). Pairwise comparisons using Tukey’s HSD showed that SI was significantly higher with medium centrality than low ($p=3.18\times 10^{-7}$), high centrality than medium ($p=1.07\times 10^{-12}$), and high centrality than low ($p=3.82\times 10^{-20}$).
@@
Recipe Complexity. Finally, we found a significant effect of recipe complexity (the number of ingredients required to make a soup) on SI. Agents tended to specialize less when recipes had one ingredient ($M=0.54,SD=0.47$) than with two ($M=0.73,SD=0.35$) or three ($M=0.67,SD=0.32$). A one-way ANOVA showed a significant difference in SI across recipes ($F(2,3097)=57.22,p=3.95\times 10^{-25}$). Pairwise comparisons using Tukey’s HSD revealed that SI was significantly higher for recipes with two ingredients compared to one ($p=4.46\times 10^{-19}$) and recipes with three ingredients compared to one ($p=4.37\times 10^{-10}$), but significantly lower for recipes with three ingredients compared to two ($p=7.77\times 10^{-6}$).
Appendix H Experiment 3: Evaluation
@@
We also fit a logistic regression model where the binary outcome variable was low versus high specialization (SI $\geq 0.5$) and the predictor variable was task parallelizability. To address the class imbalance in the SI distribution, we used class-weight balancing to automatically adjust class weights inversely proportional to their frequencies in the training data.
Figure 7: Overcooked Spatial Layouts in Experiment 3.
Figure 8: Effect of grid size on observed specialization (SI) in Experiment 3. As the size of the grid increased, agends tended to specialize more (even when presented with the same number of workstations and bottlenecks).
Spatial Layout Mean SI Mean Reward
@@
Table 3: Average specialization (SI) observed in different spatial layouts in Experiment 3.
Workstations Mean SI Std SI
@@
Table 4: Mean and standard deviation of observed specialization (SI) for varying numbers of workstations in Experiment 3.
Recipes Mean SI Std SI
@@
Table 5: Mean and standard deviation observed specialization (SI) for different recipes in Experiment 3.
@@
Table 7: Correlations between predicted task parallelizability ($S$) and observed specialization (SI) across varying spatial layouts in Experiment 3. The significance of this relationship varied across layouts, especially those with different grid sizes, prompting the controlled design of Experiments 2 and 3.
@@
Figure 9: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution grouped by a specific spatial layout.
@@
Figure 10: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution for a specific number of workstations.
@@
Figure 11: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution grouped by a specific recipe.
@@
Figure 12: Specialization (SI) distributions across all random seeds in Experiment 3. Each subfigure represents the SI distribution grouped by a specific layout. The distributions of strategies appear mostly consistent across different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
@@
Figure 13: Specialization (SI) distributions across all random seeds in Experiment 3. Each subfigure represents the SI distribution grouped by the number of workstations. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
@@
Figure 14: Specialization (SI) distributions across all random seeds in Experiment 3. Each subfigure represents the SI distribution grouped by a specific recipe. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
Appendix I Experiment 4: Results
Layout Experiment Recipe Mean SI Std SI Count Mean Reward Std Reward Reward Count
@@ Appendix I Experiment 4: Results
Table 8: Detailed results from Experiment 4. Each row represents a unique combination of spatial layout, experiment, and recipe, showing the mean, standard deviation (std), and count (number of runs) for both specialization (SI) and reward metrics. Experiment A corresponds to expanding exploration, while Experiment B denotes constant exploration. The SI count is less than 10 for certain permutations because the policy failed to converge, resulting in the agents taking no meaningful actions.
@@
Table 9: Correlations from Experiment 4. Correlations between specialization (SI) and reward for one-onion and three-onion Soup recipes under expanding and constant exploration conditions. The table reports correlations and p-values ($p$) for SI and reward in relation to layout size.
@@
Table 10: Results from Experiment 4. Pairwise T-tests comparing the smallest (Layout 1) layout with the largest layout (Layout 67).